answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many even-numbered cards in a deck of playing cards?

There are twenty even-numbered card in a standard deck of 52 playing cards, the 2, 4, 6, 8, and 10 of Spades, Diamonds, Clubs, and Hearts.

What is the probability that there are 53 Wednesdays in a leap year?

The total number of days in a leap year is 366. Then, if we want to determine the probability of 53 Wednesdays occurring in a leap year, we write 53 / 366.

What is the probabilty if a coin is tossed three times?

The probability of the coin coming up heads each time is 1/8; likewise for 3 tails. The probability of getting 2 heads and 1 tail (in any order) or 2 tails and 1 head, is 3/8. There are lots of other events whose probability can be calculated when a coin is tossed 3 times, but the question doesn't specify what event is to have its probability calculated.

How do you read a tree diagram to determine the possible outcomes for that situation?

Each leaf of the tree is one outcome. To take a simple example, consider the tree for three flips of a coin. The root splits into heads/tails, then each of the two branches splits into heads/tails, and finally each of the four branches splits into heads/tails, for 8 possible outcomes. In this simple case, each of the outcomes is equally probable. To find how many outcomes have exactly two heads, follow each branch of the tree, and put a star by the leaf each time you count two heads. You will find three leaves with exactly two heads (HHT, HTH, and THH). Add the probabilities of these three leaves, to discover that the probability of exactly three heads is 3/8.

What is the probability of choosing green or yellow?

You aren't close to providing enough information.

What is the number of possible outcomes when a coin is tossed twice?

4 : Each coin toss multiplies the generated outcome by two. The outcome for two tosses can be Heads (H) Tails (T), H H, T T or T H. As such, with three tosses the outcome can be H H H, H H T, H T H, T H H, T T H, T T T, T H T, H T T.

We can define the number of outcomes as 2^n, where n is the amount of tosses.

What is the probability of getting heads 7 times in 12 tosses of a coin?

99/512, or 19.34%.

The nCr formula can be used in this case:

12!/((12-7!)*7!) ---> 95,040/120 ---> 792

792/(2^12) = 99/512

How do I calculate probabilities for a large number of independent events without using a tree. For example if toss a coin a 1000 times how do I calculate that it will be heads only 200 times?

If p is the probability of the event happening (once), so that q = 1 - p is the probability of the event not happening (once), then the binomial theorem can be used:

(p + q)n = nC0 p0qn + nC1 p1qn-1 + ... + nCr prqn-r + ... + nCn-1 pn-1q1 + nCn pnq0

Where n is the number of trials and r is the number of successes required, and nCr is the number of combinations of r things from a set of n and is:

nCr = n!/(n-r)!r!

where n! is n factorial which is n x (n-1) x (n-2) x ... x 2 x 1. 0! is defined to be 1.

So assuming a normal, unbiased coin with p = pr(head) = 1/2 → q = pr(not_head) = 1 - 1/2 = 1/2, then the probability of getting (r =) 200 heads out of (n =) 1000 trials is:

nCr prqn-r = 1000C200 (1/2)200(1/2)800

= 1000!/(800! x 200!) (1/2)1000

How many different sum can you have when you roll 3 six sided dice?

Assuming they are standard dice numbers 1-6:

The minimum sum is 1 + 1 + 1 = 3

The maximum sum is 6 + 6 + = 18

So there are 18 - 3 + 1 = 16 different possible sums.

But note, that some of the sums may be obtained in more than one way, for example 4 can be obtained in 3 different ways: 1 + 1 + 2, 1 + 2 + 1, 2 + 1 + 1.

What do Independent trials in relationship to probability theory mean?

i dont know but why dont you do some research you lazy person

What is the probability of getting an 8 on a roll of 2 dice?

the first roll must be anything except a 1 .. odds = 5/6

the second roll must be one specific number that adds up to 8

so its odds are = 1/6

SO . . the total odds are 5/6 * 1/6 = 5/36 = about a 14% chance.

Another way to look at it is this:

there are only 5-combinatiosn that make 8 = 2/6, 3/5, 4/4, 5/3, 6/2

and there are 36 possible combinations . . .

so, again .. the odds are 5- out of 36 rolls.

Why is probability important to a scientist?

Some but not all scientific models are based on the ability to determine the likelihood that a given experimental outcome has happened by chance alone. If you have an accurate understanding of how the variables in the experiment change when nothing in particular is affecting them, then you have a way to establish some confidence that your outcome is the result of your experimental procedure and not the result of purely random events. The experimental 'lingo' is that the researcher has to determine if the 'Null Hypothesis' can be rejected. The Null Hypothesis is that the experimental outcome is not significantly different from what you would expect if the experiment had no effect at all.

As an example, if the probability in the natural world is that some event will happen by chance only one tenth of one percent of the time, then when I observe that event as my experimental outcome, I can be reasonably sure that my procedure has brought about the event; it is so unlikely that it happened by chance. It is not perfect, but few scientific procedures are. This also highlights the importance of replicating studies or of doing meta-analyses of experimental data gathered in many experiments to further reduce the likelihood that observed outcomes are nothing more than chance events.

What is probability that 2 persons sharing same birth date?

If you assume that birth dates are uniformly distributed over the year (they are not), and you ignore leap years, then the probability of two people selected at random, share a birthday is 1/365.

What is the probability of rolling a 5 on one dice 2 times?

The probability would be that of independently rolling a five once, multiplied by the same probability for a second time.

The two rolls are treated independently with P(5) = probability of rolling a single '5' = 1/6.

P(5,5) = Probability of rolling one 5 and then another 5, out of two rolls

p(5,5)=p(5 AND 5)=p(5)*p(5)=(1/6)*(1/6)=1/36

So it's 1/36, or 0.0278

When a two dice is rolled what is the probability the two number will have a sum of 8?

5/36

Out of the 36 possible and equally-probable outcomes (1st die's number and 2nd die's number), 5 combinations give you a sum equal 8:

* 1st die gives a 2 and 2nd die gives a 6

* 1st die gives a 6 and 2nd die gives a 2

* 1st die gives a 3 and 2nd die gives a 5

* 1st die gives a 5 and 2nd die gives a 3

* 1st die gives a 4 and 2nd die gives a 4

Therefore the probability that the sum is greater than 10 equals 5/36.

How many zeros are possible factorial 10 to the power factorial 10?

To calculate the number of zeros in a factorial number, we need to determine the number of factors of 5 in the factorial. In this case, we are looking at 10 to the power of 10 factorial. The number of factors of 5 in 10! is 2 (from 5 and 10). Therefore, the number of zeros in 10 to the power of 10 factorial would be 2.

How would you explain the concept of probability to a child?

Start with examples like flipping a coin, rolling a die or spinning a dreidel. Then explain in terms they understand. That depends very much on the age of the child.

How the probability and statistics is apply in insurance?

They are used to understand how to set insurance premiums.

Take, for example, car insurance.

The probability of a new driver getting in an accident is higher than an experienced driver.

The probability of a male driver getting in an accident is higher than an female driver (in general).

The probability of a person who has had many claims filing another claim is higher than someone who has never filed a claim.

All these data guide the insurance industry to charge higher premiums to drivers who are new, male or have filed claims in the past.